# Amortized Active Generation of Pareto Sets

### v2

@@ 6 Experiments
We now evaluate A-GPS on a number of benchmarks and compare it to some popular baselines. Firstly we apply A-GPS to a number of well known continuous synthetic MOO test functions as a proof-of-concept. Then we apply it to three high-dimensional sequence design challenges — A-GPS’s intended application — that emulate real protein engineering tasks. Our primary measure of performance is Pareto front relative HVI [42, 16] using the implementation in [3]. For all experiments we set $\beta=0.5$ as the full KL regularization in Equation 15 can hamper exploitation in later rounds on some tasks. We refer the reader to Appendix D for full experimental details.
@@ 6.1 Synthetic test functions
As a proof-of-concept, we demonstrate A-GPS on some classical continuous paramterized MOO problems ($\mathbf{x}\in\real^{D}$) commonly used for MOBO [41, 4, 3]. Even though A-GPS has not been designed for purely continuous problems, they none-the-less allow us to demonstrate some appealing properties of A-GPS. We present three here; negative Branin-Currin ($D=2$, $L=2$), DTLZ7 ($D=7$, $L=6$), and ZDT3 ($D=4$, $L=2$), see [13, 41, 4]. More detailed descriptions of these functions, additional experimental detail and on additional test functions are presented in Sec. D.1. We apply A-GPS to these continuous domains using a conditional Gaussian generative model, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})=\mathcal{N}\!\left({\mathbf{x}}\middle|{\boldsymbol{\mu}(\mathbf{u}),\boldsymbol{\sigma}^{2}(\mathbf{u})}\right)$, with mean and variance parameterized by a neural network (NN). The top row of Figure 2 reports mean relative hypervolume versus optimization round. The bands indicating $\pm 1$ std. from 10 runs with random parameter initialisation. We compare to three Gaussian process (GP)-based baselines, qNEHVI [8], qEHVI and qNParEGO [7]. All methods use 64 training points, and then recommend $B=5$ candidates for $T=10$ rounds, and A-GPS has $\tau_{0}$ set using the $p=0.25$ percentile of Pareto ranks. On Branin-Currin, A-GPS (purple) rapidly outpaces the scalarization based qNParEGO, but is dominated by the methods that explicitly estimate EHVI. However, these methods do not scale to the higher dimensional DTLZ7 problem, where A-GPS still outperforms qNParEGO. While A-GPS can model the complex Pareto front of ZDT3, the direct optimization methods outperform it, we suspect a stronger generative backbone for continuous data would help here.
@@
We now evaluate A-GPS on a challenging two-objective synthetic ‘peptide’ design task that couples the Ehrlich synthetic landscape [33] with a ProtBert [15] ‘naturalness’ score. The Ehrlich function has been designed to emulate key aspects of protein fitness; it maps each discrete sequence to a scalar by embedding combinatorial motif interactions in a highly rugged, multi-modal, but artificial terrain. In stark contrast, ProtBert’s loss reflects genuine amino-acid patterns learned from 217 million real proteins. We use protein sequences $\mathcal{X}=\mathcal{V}^{M}$ where $|\mathcal{V}|=20$ and $M\in\{15,32,64\}$. The two objectives are,
@@
Here we convert ProtBert’s log-loss, $\mathcal{L}_{\text{ProtBert}}(\mathbf{x})$ into a likelihood (${f\!\centerdot}^{2}\in[0,1]$), making it output in a comparable range to the Ehrlich function (${f\!\centerdot}^{1}\in\{-1\}\cup[0,1]$). We compare against a baseline that randomly mutates the set of best candidates, CbAS [5] and VSD [35], which use the same 2-layer CNN Pareto CPE as A-GPS, and against the guided diffusion based LaMBO-2 [18], which is formulated for discrete MOBO tasks using EHVI guidance. A-GPS uses the same CNN architecture for its alignment CPE. CbAS and VSD use a causal transformer architecture for their $q_{\phi}({\mathbf{x}})$. A-GPS uses the same backbone transformer, but embeds $\mathbf{u}$ for its prefix token and uses FiLM [29] on the transformer embeddings for conditioning the transformer on preferences, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$. The same (unconditional) architectures are used as priors by CbAS, VSD and A-GPS, which are trained on the initial sequences using maximum likelihood. Unlike in [35, 5], we allow CbAS to resample $q_{\phi}({\mathbf{x}})$ between rounds (once every 100 iterations) — this drastically improves its performance. Results are reported in Figure 3 for $T=40$ rounds from random starting conditions (bands indicating $\pm 1$ std). All methods are given 128 training samples, and then recommend batches of size $B=32$ per round, and $\tau_{0}$ is set as $p=0.25$ percentile of Pareto ranks. We use the poli and poli-baselines libraries for running the benchmarks and LaMBO-2 baseline [17]. Additional experimental details and timing results are in Sec. D.2. A-GPS performs similarly to VSD, CbAS tends to be overly exploitative, and LaMBO-2’s guided masked-diffusion model tends to under-perform on this task.
@@ 6.3 Bi-grams
For this experiment we use the bi-grams optimization task from [34] where the aim is to maximize the occurance of three bi-grams (‘AV’, ‘VC’ and ‘CA’) in an $M=32$ length sequence. We start with 512 random sequences that have no more than three of these bi-grams present. We then have $T=64$ rounds of $B=16$ to optimize the bi-gram occurances. The initial Pareto front is sparse so $\tau_{0}$ is the $p=0.5$ percentile of Pareto ranks. The same models as the previous experiment are used along with a masked-transformer model (mTFM) backbone for CbAS, VSD and A-GPS. This allows control over the number of mutations to apply to an existing sequence, rather than generating a complete sequence from scratch — see Appendix C for details. LaMBO-2’s masked diffusion backbone also has this ability, and following [34] we use a 1-mutation budget for these models. Relative hypervolume improvement results are presented in Figure 4 (a) and (b), sequence diversity is computed as the average pair-wise edit distance between all sequences. We can see the causal transformer backbone A-GPS and VSD models perform best, followed by LaMBO-2. CbAS overfits early as we can see from the diversity plot, and the mTFM models perform similarly to the random baseline and take many rounds to show any improvement on this task.
@@
Our final experiment uses the simulation-based protein stability vs. solvent accessible surface area (SASA) task from [34]. The aim is to optimize six base red fluorescent proteins with $M>200$ for stability ($\Delta G$) and SASA. We use the FoldX black-box implementation in poli [17], with 512 training samples, $T=64$, $B=16$, and a budget of one mutation per round. We have a rich starting Pareto front so $\tau_{0}$ is set from $p=0.1$. We only use the mTFM backbone for CbAS, VSD and A-GPS as FoldX is best modelling only small differences to the original sequences. The results are summarised in Figure 4 (c) and (d). LaMBO-2 initially performs well, but is overcome by A-GPS as the sequence diversity diminishes. See Sec. D.4 for additional plots of the estimated Pareto front.
@@ 7 Limitations and Discussion
A limitation with A-GPS, and one that it shares with many MOBO and MOG methods, is that it can be hard to specify algorithm hyper-parameters a-priori — before new data has been acquired — and the settings of these hyper-parameters can effect real-world performance. We are mindful of this in our implementation and design of A-GPS, and as such it comprises components that can be independently trained and validated meaningfully on the initial training data at hand. In particular, we find A-GPS, VSD and CbAS are sensitive to the prior model used, $p({\mathbf{x}}|{\mathcal{D}_{0}})$. To aid practitioners use these methods, we outline a general procedure we find works well in Sec. B.2 for prior choice and initialization. Another limitation with A-GPS as presented in this work is the choice of the generative model for the continuous synthetic test functions. The MLP used does not appear to scale well to higher-dimensional problems, and we expect using flow-matching [26] or diffusion [20] backbones would remedy this issue, and incorporating these models into the VSD and A-GPS frameworks is a future research direction.
@@ Acknowledgements
This work is funded by the CSIRO Science Digital and Advanced Engineering Biology Future Science Platforms, and was supported by resources and expertise provided by CSIRO IMT Scientific Computing. We would like to thank the anonymous reviewers for the constructive feedback and advice, which greatly increased the quality of this work.
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. 2.
    Limitations
  7. Answer: [Yes]
  9. 3.
    Theory assumptions and proofs
  11. Answer: [Yes]
  13. 4.
    Experimental result reproducibility
  15. Answer: [Yes]
  17. 5.
    Open access to data and code
  19. Answer: [Yes]
  21. 6.
    Experimental setting/details
  23. Answer: [Yes]
  25. 7.
    Experiment statistical significance
  27. Answer: [Yes]
  29. 8.
    Experiments compute resources
  31. Answer: [Yes]
  33. 9.
    Code of ethics
  35. Answer: [Yes]
  37. 10.
    Broader impacts
  39. Answer: [Yes]
  41. 11.
    Safeguards
  43. Answer: [N/A]
  45. 12.
    Licenses for existing assets
  47. Answer: [Yes]
  49. 13.
    New assets
  51. Answer: [Yes]
  53. 14.
    Crowdsourcing and research with human subjects
  55. Answer: [N/A]
  57. 15.
  59. Answer: [N/A]
  61. 16.
    Declaration of LLM usage
  63. Answer: [Yes]
@@ B.1 A Non-dominance CPE is estimating probability of hypervolume improvement
We define a box of hypervolume for any $\mathbf{y}\in\mathcal{Y}$ under 1 as,
@@
Case 2: $\mathbf{x}$ is non-dominated. If $\mathbf{x}^{\prime}\not\succ\mathbf{x}$ for all $\mathbf{x}^{\prime}\in\mathcal{S}_{\textrm{Pareto}}^{t}$, there is no $\mathbf{y}^{\prime}\in\mathcal{F}_{\textrm{Pareto}}^{t}$ such that $\mathrm{B}({\mathbf{y}})\subset\mathrm{B}({\mathbf{y}^{\prime}})$. For $\mathrm{HVI}({\mathbf{x}})$ to be positive, we need to show that $\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}})<\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}\cup\{\mathbf{y}\}})$, i.e., whatever remains of the difference between $\mathrm{B}({\mathbf{y}})$ and the previous union $\bigcup_{\mathbf{y}^{\prime}\in\mathcal{F}_{\textrm{Pareto}}^{t}}\!\mathrm{B}({\mathbf{y}^{\prime}})$ must be a set of positive Lebesgue measure. Indeed, letting $\gamma:=\min_{\mathbf{x}^{\prime}\in\mathcal{S}_{\textrm{Pareto}}^{t}}\max_{i\in\{1,\dots,L\}}y_{i}-y^{\prime}_{i}$, which is positive due to dominance, and setting $\delta:=\min\{\gamma/2,\min_{i\in\{1,\dots,L\}}(y_{i}-r_{i})/2\}>0$ (by 1), we have that the $\delta$-box $\mathrm{B}_{\delta}:=\prod_{i=1}^{L}[y_{i}-\delta,y_{i}]\subset\mathrm{B}({\mathbf{y}})$ is not covered by the union $\bigcup_{\mathbf{y}^{\prime}\in\mathcal{F}_{\textrm{Pareto}}^{t}}\!\mathrm{B}({\mathbf{y}^{\prime}})$. It then follows that $\mathrm{HVI}({\mathbf{x}})=\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}\cup\{\mathbf{y}\}})-\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}})\geq\lambda_{L}(\mathrm{B}_{\delta})>0$, confirming that
@@ B.2 Fitting the prior
    Train prior with 10% validation set, make note of number of iterations when validation loss begins to increase
@@ C.1 Preference direction distributions
and $\gamma=\{(\boldsymbol{\mu}_{k},\boldsymbol{\sigma}_{k})\}^{K}_{k=1}$. Typically, we find $K=5$ is sufficient. We learn this via maximum likelihood as per Equation 16, but we add an extra regularisation term: $\frac{1}{K}\sum^{K}_{k=1}(\|\boldsymbol{\mu}_{k}\|-1)^{2}$ so the magnitude of the mixture means is controlled (and does not decrease to 0 or increase to $\pm\infty$). We have compared this to von Mises distributions, and find it more numerically stable, we also find no tangible benefit using more complex spherical normalizing flow representations [32]. Furthermore, we find that the performance is similar to, if not slightly superior to, the empirical approximation,

### v3

@@ 6 Experiments
We now evaluate A-GPS on a number of benchmarks and compare it to some popular baselines. Firstly we apply A-GPS to a number of well known continuous synthetic MOO test functions as a proof-of-concept. Then we apply it to three high-dimensional sequence design challenges — A-GPS’s intended application — that emulate real protein engineering tasks. Our primary measure of performance is Pareto front relative HVI [zitzler1999multiobjective, fonseca2006improved] using the implementation in [balandat2020botorch]. For all experiments we set $\beta=0.5$ as the full KL regularization in Equation 15 can hamper exploitation in later rounds on some tasks. We refer the reader to Appendix D for full experimental details.
@@ 6.1 Synthetic test functions
As a proof-of-concept, we demonstrate A-GPS on some classical continuous paramterized MOO problems ($\mathbf{x}\in\real^{D}$) commonly used for MOBO [zhang2007moea, belakaria2019max, balandat2020botorch]. Even though A-GPS has not been designed for purely continuous problems, they none-the-less allow us to demonstrate some appealing properties of A-GPS. We present three here; negative Branin-Currin ($D=2$, $L=2$), DTLZ7 ($D=7$, $L=6$), and ZDT3 ($D=4$, $L=2$), see [deb2002scalable, zhang2007moea, belakaria2019max]. More detailed descriptions of these functions, additional experimental detail and on additional test functions are presented in Sec. D.1. We apply A-GPS to these continuous domains using a conditional Gaussian generative model, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})=\mathcal{N}\!\left({\mathbf{x}}\middle|{\boldsymbol{\mu}(\mathbf{u}),\boldsymbol{\sigma}^{2}(\mathbf{u})}\right)$, with mean and variance parameterized by a neural network (NN). The top row of Figure 2 reports mean relative hypervolume versus optimization round. The bands indicating $\pm 1$ std. from 10 runs with random parameter initialisation. We compare to three Gaussian process (GP)-based baselines, qNEHVI [daulton2021parallel], qEHVI and qNParEGO [daulton2020differentiable]. All methods use 64 training points, and then recommend $B=5$ candidates for $T=10$ rounds, and A-GPS has $\tau_{0}$ set using the $p=0.25$ percentile of Pareto ranks. On Branin-Currin, A-GPS (purple) rapidly outpaces the scalarization based qNParEGO, but is dominated by the methods that explicitly estimate EHVI. However, these methods do not scale to the higher dimensional DTLZ7 problem, where A-GPS still outperforms qNParEGO. While A-GPS can model the complex Pareto front of ZDT3, the direct optimization methods outperform it, we suspect a stronger generative backbone for continuous data would help here.
@@
We now evaluate A-GPS on a challenging two-objective synthetic ‘peptide’ design task that couples the Ehrlich synthetic landscape [stanton2024closed] with a ProtBert [elnaggar2020prottrans] ‘naturalness’ score. The Ehrlich function has been designed to emulate key aspects of protein fitness; it maps each discrete sequence to a scalar by embedding combinatorial motif interactions in a highly rugged, multi-modal, but artificial terrain. In stark contrast, ProtBert’s loss reflects genuine amino-acid patterns learned from 217 million real proteins. We use protein sequences $\mathcal{X}=\mathcal{V}^{M}$ where $|\mathcal{V}|=20$ and $M\in\{15,32,64\}$. The two objectives are,
@@
Here we convert ProtBert’s log-loss, $\mathcal{L}_{\text{ProtBert}}(\mathbf{x})$ into a likelihood (${f\!\centerdot}^{2}\in[0,1]$), making it output in a comparable range to the Ehrlich function (${f\!\centerdot}^{1}\in\{-1\}\cup[0,1]$). We compare against a baseline that randomly mutates the set of best candidates, CbAS [brookes2019conditioning] and VSD [steinberg2025variational], which use the same 2-layer CNN Pareto CPE as A-GPS, and against the guided diffusion based LaMBO-2 [gruver2023protein], which is formulated for discrete MOBO tasks using EHVI guidance. A-GPS uses the same CNN architecture for its alignment CPE. CbAS and VSD use a causal transformer architecture for their $q_{\phi}({\mathbf{x}})$. A-GPS uses the same backbone transformer, but embeds $\mathbf{u}$ for its prefix token and uses FiLM [perez2018film] on the transformer embeddings for conditioning the transformer on preferences, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$. The same (unconditional) architectures are used as priors by CbAS, VSD and A-GPS, which are trained on the initial sequences using maximum likelihood. Unlike in [steinberg2025variational, brookes2019conditioning], we allow CbAS to resample $q_{\phi}({\mathbf{x}})$ between rounds (once every 100 iterations) — this drastically improves its performance. Results are reported in Figure 3 for $T=40$ rounds from random starting conditions (bands indicating $\pm 1$ std). All methods are given 128 training samples, and then recommend batches of size $B=32$ per round, and $\tau_{0}$ is set as $p=0.25$ percentile of Pareto ranks. We use the poli and poli-baselines libraries for running the benchmarks and LaMBO-2 baseline [gonzalez-duque2024poli]. Additional experimental details and timing results are in Sec. D.2. A-GPS performs similarly to VSD, CbAS tends to be overly exploitative, and LaMBO-2’s guided masked-diffusion model tends to under-perform on this task.
@@ 6.3 Bi-grams
For this experiment we use the bi-grams optimization task from [stanton2022accelerating] where the aim is to maximize the occurance of three bi-grams (‘AV’, ‘VC’ and ‘CA’) in an $M=32$ length sequence. We start with 512 random sequences that have no more than three of these bi-grams present. We then have $T=64$ rounds of $B=16$ to optimize the bi-gram occurances. The initial Pareto front is sparse so $\tau_{0}$ is the $p=0.5$ percentile of Pareto ranks. The same models as the previous experiment are used along with a masked-transformer model (mTFM) backbone for CbAS, VSD and A-GPS. This allows control over the number of mutations to apply to an existing sequence, rather than generating a complete sequence from scratch — see Appendix C for details. LaMBO-2’s masked diffusion backbone also has this ability, and following [stanton2022accelerating] we use a 1-mutation budget for these models. Relative hypervolume improvement results are presented in Figure 4 (a) and (b), sequence diversity is computed as the average pair-wise edit distance between all sequences. We can see the causal transformer backbone A-GPS and VSD models perform best, followed by LaMBO-2. CbAS overfits early as we can see from the diversity plot, and the mTFM models perform similarly to the random baseline and take many rounds to show any improvement on this task.
@@
Our final experiment uses the simulation-based protein stability vs. solvent accessible surface area (SASA) task from [stanton2022accelerating]. The aim is to optimize six base red fluorescent proteins with $M>200$ for stability ($\Delta G$) and SASA. We use the FoldX black-box implementation in poli [gonzalez-duque2024poli], with 512 training samples, $T=64$, $B=16$, and a budget of one mutation per round. We have a rich starting Pareto front so $\tau_{0}$ is set from $p=0.1$. We only use the mTFM backbone for CbAS, VSD and A-GPS as FoldX is best modelling only small differences to the original sequences. The results are summarised in Figure 4 (c) and (d). LaMBO-2 initially performs well, but is overcome by A-GPS as the sequence diversity diminishes. See Sec. D.4 for additional plots of the estimated Pareto front.
@@ 7 Limitations and Discussion
A limitation with A-GPS, and one that it shares with many MOBO and MOG methods, is that it can be hard to specify algorithm hyper-parameters a-priori — before new data has been acquired — and the settings of these hyper-parameters can effect real-world performance. We are mindful of this in our implementation and design of A-GPS, and as such it comprises components that can be independently trained and validated meaningfully on the initial training data at hand. In particular, we find A-GPS, VSD and CbAS are sensitive to the prior model used, $p({\mathbf{x}}|{\mathcal{D}_{0}})$. To aid practitioners use these methods, we outline a general procedure we find works well in Sec. B.2 for prior choice and initialization. Another limitation with A-GPS as presented in this work is the choice of the generative model for the continuous synthetic test functions. The MLP used does not appear to scale well to higher-dimensional problems, and we expect using flow-matching [lipman2022flow] or diffusion [ho2020denoising] backbones would remedy this issue, and incorporating these models into the VSD and A-GPS frameworks is an active research direction.
Another avenue of future work would be to explore the rich literature on MOEAs for alternative multi-objective measures beyond scalarizations and hypervolume indicators that may be suitable for guided MOG [afsar2023many]. Additionally, this literature gives a thorough treatment of stochastic experiments in the multi-objective setting [rojasgonzalez2025bi] which may be useful for MOG.
@@ Acknowledgements
This work is funded by the CSIRO Science Digital and Advanced Engineering Biology Future Science Platforms, and was supported by resources and expertise provided by CSIRO IMT Scientific Computing. We would like to thank the anonymous reviewers for the constructive feedback and advice, and Sebastian Rojas Gonzalez for pointing us to the literature on MOEAs and for the discussion on estimating PHVI— all of which greatly increased the quality of this work.
@@ B.1 A Non-dominance CPE is estimating probability of hypervolume improvement
Assumption 1 (Measurement noise).
No objectives exhibit measurement noise, $\boldsymbol{\epsilon}=\mathbf{0}$.
This is so our dominance comparison, $\succ$, and hypervolume computation are exact with no ambiguity introduced from noisy measurements. We will discuss the consequences of relaxing this assumption later.
@@
We define a box of hypervolume for any $\mathbf{y}\in\mathcal{Y}$ under 1 and 2 as,
@@
Case 2: $\mathbf{x}$ is non-dominated. If $\mathbf{x}^{\prime}\not\succ\mathbf{x}$ for all $\mathbf{x}^{\prime}\in\mathcal{S}_{\textrm{Pareto}}^{t}$, there is no $\mathbf{y}^{\prime}\in\mathcal{F}_{\textrm{Pareto}}^{t}$ such that $\mathrm{B}({\mathbf{y}})\subset\mathrm{B}({\mathbf{y}^{\prime}})$. For $\mathrm{HVI}({\mathbf{x}})$ to be positive, we need to show that $\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}})<\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}\cup\{\mathbf{y}\}})$, i.e., whatever remains of the difference between $\mathrm{B}({\mathbf{y}})$ and the previous union $\bigcup_{\mathbf{y}^{\prime}\in\mathcal{F}_{\textrm{Pareto}}^{t}}\!\mathrm{B}({\mathbf{y}^{\prime}})$ must be a set of positive Lebesgue measure. Indeed, letting $\gamma:=\min_{\mathbf{x}^{\prime}\in\mathcal{S}_{\textrm{Pareto}}^{t}}\max_{i\in\{1,\dots,L\}}y_{i}-y^{\prime}_{i}$, which is positive due to dominance, and setting $\delta:=\min\{\gamma/2,\min_{i\in\{1,\dots,L\}}(y_{i}-r_{i})/2\}>0$ (by 2), we have that the $\delta$-box $\mathrm{B}_{\delta}:=\prod_{i=1}^{L}[y_{i}-\delta,y_{i}]\subset\mathrm{B}({\mathbf{y}})$ is not covered by the union $\bigcup_{\mathbf{y}^{\prime}\in\mathcal{F}_{\textrm{Pareto}}^{t}}\!\mathrm{B}({\mathbf{y}^{\prime}})$. It then follows that $\mathrm{HVI}({\mathbf{x}})=\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}\cup\{\mathbf{y}\}})-\mathrm{HV}({\mathcal{F}_{\textrm{Pareto}}^{t}})\geq\lambda_{L}(\mathrm{B}_{\delta})>0$, confirming that
@@
It is worth noting that we do not require 1 for these indicators to be equivalent so long as the same noisy samples, $\mathbf{y}_{i}$, are being used when computing them. However, we can no longer claim that the indicators themselves are reliably computing hypervolume improvement or non-dominance since the addition of noise makes the underlying comparisons ambiguous with respect to the true objectives, $\mathbf{f}\centerdot$}. It may still be possible to show convergence of A-GPS to the true Pareto set, $\mathcal{S}_{\textrm{Pareto}}$, while using noisy observations. The authors of VSD [steinberg2025variational] were able to guarantee convergence using a CPE trained with noisy PI labels for single objective BBO using results from the neural tangent kernel (NTK) literature (see their Appendix F). Some of these results may be extended to Pareto set classification, though we leave this for future work, as all of our experiments are noise-free.
@@ B.2 Fitting the prior
    Train prior with a 10% validation set, make note of number of iterations when validation loss begins to increase
@@ C.1 Preference direction distributions
and $\gamma=\{(\boldsymbol{\mu}_{k},\boldsymbol{\sigma}_{k})\}^{K}_{k=1}$. Typically, we find $K=5$ is sufficient. We learn this via maximum likelihood as per Equation 16, but we add an extra regularisation term: $\frac{1}{K}\sum^{K}_{k=1}(\|\boldsymbol{\mu}_{k}\|-1)^{2}$ so the magnitude of the mixture means is controlled (and does not decrease to 0 or increase to $\pm\infty$). We have compared this to von Mises distributions, and find it more numerically stable, we also find no tangible benefit using more complex spherical normalizing flow representations [rezende2020normalizing]. Furthermore, we find that the performance is similar to, if not slightly superior to, the empirical approximation,